y=-1/4x-5, (7,-5) write an equation of the line in slope intercept from that is perpendicular to the given line and passes through the given point.
Answer by algebrahouse.com(1659)

"y = -1/4x - 5, (7,-5) write an equation of the line in slope intercept from that is perpendicular to the given line and passes through the given point."

Perpendicular lines have slopes which are negative reciprocals of each other.

In the equation, y = -1/4x - 5
slope = -1/4

The slope of the line perpendicular to that, would be 4 {negative reciprocal of -1/4}

(7,-5), m = 4 {the point and slope of the desired line}
y = mx + b {slope-intercept form for the equation of a line}
-5 = 4(7) + b {substituted into slope-intercept form}
-5 = 28 + b {multiplied 4 by 7}
b = 33 {subtracted 28 from both sides}
- the y-intercept is 33

m = 4, b = 33 {slope and y-intercept}
y = mx + b {slope-intercept form}
y = 4x + 33 {substituted into slope-intercept form}

the equation of the line is y = 4x + 33
www.algebrahouse.com